# Map enumeration from a dynamical perspective

### v2

@@ Abstract.
Mathematics Subject Classifications: 37J65, 37J70; Secondary: 39A30, 39A60, 41A60
${}^{1}$ University of Arizona, Department of Mathematics (ercolani@math.arizona.edu)
${}^{2}$ University of Arizona, Department of Mathematics (lega@math.arizona.edu, www.math.arizona.edu/~lega/),
${}^{3}$ University of Arizona, Department of Mathematics (tippings@arizona.edu).
@@ 1. Introduction: map enumeration, random matrices, and orthogonal polynomials
Given a compact oriented surface of genus $g$, enumerating maps means finding how many different maps with $j$ vertices can be embedded in the surface. To address the issue of permuting vertices, a common approach is to label them ahead of time. Accounting for all of the vertex labelings therefore introduces a factor of $j!$ in the map count. To describe the graph, one also needs to label the darts, which are the half-edges emanating from each vertex. Pairing two darts defines an edge between the vertices to which they are attached. Because the surface is oriented and the vertices are already numbered, the order in which the dart labels are arranged about each vertex is fixed. In the case of a regular-valence map, such that the degree of each vertex is $2\nu$, labeling the darts therefore results in an additional factor of $(2\nu)^{j}$, where the $2\nu$ comes from rotating the dart numbers counterclockwise about each vertex. When enumerating maps, one then needs to specify whether counts are for labeled maps or unlabeled maps, the former being $(2\nu)^{j}\cdot j!$ as large as the latter. Another consideration that affects map counts is related to symmetries of the surface. Indeed, two maps are considered equivalent (i.e. not different) if there is an orientation-preserving homeomorphism of the surface onto itself that leaves the map unchanged. This may lead to unlabeled counts that are rational numbers. An illustrative example is provided in Figure 1. On a surface of genus 0, there are a priori four 4-valent labeled maps with one vertex. These are shown on the left-hand side of Figure 1 and are obtained from one-another by rotating counterclockwise the labeling of the 4 darts. However, because the plane (or sphere) is invariant under rotations, the first and third maps are the same. So are the second and fourth (red double arrows). There are therefore two different labeled maps, one for each box on the right-hand-side of the figure. Consequently, the number of different unlabeled maps is $22/(4^{1}\cdot 1!)=1/$ 2, which is not integral.
@@
[…]
@@
where $z_{0}=z_{0}(t,\alpha)$ and $P_{3g-2}$ is a polynomial of degree $23g-$ 2. This is a priori surprising since $z_{g}$, which may be used to count maps with an arbitrary number of vertices, only depends on $z_{0}$ (which is explicitly known from the string equation (1.12) below) and a finite number of parameters: the half-valence $\nu$ and the $13g-$ 1 coefficients of $P_{3g-2}$. The connection with the generating functions $e_{g}$ was made in [Er14], which established that $e_{g}(t,\alpha)$ solves a Cauchy-Euler equation in $t$, with a forcing term that depends on the generating functions $\{z_{k}(t,\alpha),\ k\leq g\}$ and $\{e_{k}(t,\alpha),\ k<g\}$. Building on this result, [Er14] demonstrated that for $g\geq 2$, $e_{g}(t,\alpha)$ is also a rational function of $z_{0}=z_{0}(t,\alpha)$,
@@
where $Q_{5g-5-r}$ is a polynomial of degree $5g-5-r$. This, which proved and extended a conjecture proposed for $\nu=2$ in [BIZ80], led to the partial fraction expansion
@@
These results made it possible to express the number of regular even-valent $g$-maps with $j$ vertices in terms of $z_{0}$ and a finite number of real coefficients (specifically $33g+j-$ 3). Indeed, taking successive derivatives of (1.11) with respect to $t$ gives, for $j\geq 1$, [Er14]
@@ 2.1. Orthogonal polynomials, Painlevé equation, and center manifold
[…]
@@
[…]
@@
Since $x_{n}\to\infty$ as $n\to\infty$ for the Freud orbit, the corresponding sequence $\{u_{n}\}$ converges to 00. In addition since $x_{n}\sim y_{n}$ the sequence $\{s_{n}\}$ converges to $2$, and using (2.4) we see that the sequence $\{f_{n}\}$ converges to $2$ as well. In other words, the above transformation has the property that the Freud orbit converges to the fixed point $P_{\infty}$ with coordinates $(2,2,0)$ in the $(s,f,u)$ space.
@@ 2.3. The 4-valent case
Although the number, $33g+j-$ 3, of possibly non-zero entries of $V^{(j)}$ increases with $j$, some of these coefficients may nevertheless vanish. As shown in [ELT23b], when $\nu=2$, the dynamics of the sequence $\{V^{(j)}\}$ simplifies: for $g\geq 2$, all of the potentially non-zero entries of $V^{(j)}$ are consecutively found in a band whose length, $55g-$ 5, is independent of $j$. Consequently, the dynamics of the sequence $\{V^{(j)}\}$ can be recast into that of a finite-dimensional discrete dynamical system by writing
@@
where the $x_{k}^{(j)}$ represent the entries of $V^{(j)}$ that are possibly non-zero, and defining the $(5g-5)$-dimensional map
@@ 4.1. Derivation of the mixed valence discrete dynamical systems
[…]
@@ 4.2. Center manifold
Linearization about $P_{\infty}$ reveals that the five eigenvalues are $\lambda=1$ with multiplicity one, and $\lambda=-2\pm\sqrt{3}$, each with multiplicity two. Consequently, $P_{\infty}$ has a one-dimensional center direction. A center manifold expansion may be obtained by writing $s,\ f,\ z$, and $w$ in powers of $u$, and solving order by order after demanding that the resulting curve be invariant under the dynamics of system (4.4) and tangent to the center direction at $P_{\infty}$. Such a calculation can easily be performed with the assistance of a computer algebra software. Using Maple [Map21], we obtain the following asymptotic expansion as $u\to 0$,
@@ 5. Conclusions: open problems and challenges
Mixed 3 & 4
In progress
4.2]
In progress
In progress
@@
Valence
4
$2\nu,\ \nu\in\mathbb{N}$
3
Mixed 3 & 4
Results Related to   Map Counts
$\bullet\ 0\leq g\leq 2$ [EMP08] $\bullet\ 0\leq g\leq 7$ (Recursive) [ELT23a] $\bullet\ 1\leq g\leq 7$ (Explicit) [ELT23b]
$\bullet\ 0\leq g\leq 2$ [EMP08]    $\bullet\ 2\leq g\quad$ (Hyper-     geometric functions) [ELT23b]
$\bullet\ 0\leq g\leq 2$ [EP12]
In progress
Previously Known Results Related to    Map Counts
$\bullet\ 0\leq g\leq 2$ [BIZ80] $\bullet\ 0\leq g\leq 5$ [Pi06]        $\bullet\ 0\leq g\leq 5$ [DY17]      $\bullet\ g=3\qquad$ [BGM21]
$\bullet\ 0\leq g\leq 2$ [BD13] $\bullet\ 0\leq g\leq 5$ [DY17]
@@ 5.1.2. Center manifold expansion
[…]
@@ 5.2.1. An explicit identity involving the Gauss hypergeometric function
We have numerically verified the validity of Conjecture 5.1 for values of $\ell$ between 0 and 20, and values of $g$ and $j$ between 1 and 20.
@@ 5.2.3. Interlacing zeros of generating functions for 2-legged maps
where $S_{g-1}$ is a polynomial of degree $g-1$. In [ELT23a], we established such expressions for values of $g$ up to 7 and noticed that the zeros of the polynomials $S_{g-1}$, for $2\leq g\leq 7$, are all real and interlace. Further automation (with the help of a symbolic computation software) of the comparison of the genus and center manifold expansions performed in [ELT23a] will indicate whether this behavior continues for larger values of $g$. Using an alternate method to calculate the $z_{g}$’s, we were able to further verify this trend for $g$ up to 15. If this continues, understanding the structure of these polynomials, as well as the nature of the distribution of their zeros as $g\to\infty$, promises to be a rich and rewarding line of investigation.
@@ 5.2.4. Higher order asymptotic expansions for map counts
thereby clarifying the rate at which ${\mathcal{N}}_{2\nu,e}(g,j)$ converges to ${\mathcal{N}}_{2\nu,e}^{\infty}(g,j)$. Plots of the ratios ${\mathcal{N}}_{2\nu,e}(g,j)\over{\mathcal{N}}_{2\nu,e}^{\infty}(g,j)$ are provided in [ELT23b] for $\nu=2$ and $2\leq g\leq 7$.
@@ 5.2.5. Map counts and Bernoulli numbers
When proven, the above conjecture will provide a novel method to calculate the orbifold Euler characteristic $\chi_{orb}(M_{g})$ from the generating functions $e_{g}(z_{0})$.

### v3

@@ Abstract.
2020 Mathematics Subject Classification:
Primary: 37J65, 37J70; Secondary: 39A30, 39A60, 41A60
@@ 1. Introduction: map enumeration, random matrices, and orthogonal polynomials
Given a compact oriented surface of genus $g$, enumerating maps means finding how many different maps with $j$ vertices can be embedded in the surface. To address the issue of permuting vertices, a common approach is to label them ahead of time. Accounting for all of the vertex labelings therefore introduces a factor of $j!$ in the map count. To describe the graph, one also needs to label the darts, which are the half-edges emanating from each vertex. Pairing two darts defines an edge between the vertices to which they are attached. Because the surface is oriented and the vertices are already numbered, the order in which the dart labels are arranged about each vertex is fixed. In the case of a regular-valence map, such that the degree of each vertex is $2\nu$, labeling the darts therefore results in an additional factor of $(2\nu)^{j}$, where the $2\nu$ comes from rotating the dart numbers counterclockwise about each vertex. When enumerating maps, one then needs to specify whether counts are for labeled maps or unlabeled maps, the former being $(2\nu)^{j}\cdot j!$ as large as the latter. Another consideration that affects map counts is related to symmetries of the surface. Indeed, two maps are considered equivalent (i.e. not different) if there is an orientation-preserving homeomorphism of the surface onto itself that leaves the map unchanged. This may lead to unlabeled counts that are rational numbers. An illustrative example is provided in Figure 1. On a surface of genus 0, there are a priori four 4-valent labeled maps with one vertex. These are shown on the left-hand side of Figure 1 and are obtained from one-another by rotating counterclockwise the labeling of the 4 darts. However, because the plane (or sphere) is invariant under rotations, the first and third maps are the same. So are the second and fourth (red double arrows). There are therefore two different labeled maps, one for each box on the right-hand-side of the figure. Consequently, the number of different unlabeled maps is $2/(4^{1}\cdot 1!)=1/2$, which is not integral.
@@
$\widehat{Z}_{N}^{(n)}(\mathbf{t})=\mathop{\text{\huge$\int$}}\nolimits\cdots\mathop{\text{\huge$\int$}}\nolimits\exp\left\{-N^{2}\left[\frac{1}{N}\sum_{j=1}^{n}V_{{\bf t}}(\lambda_{j})-\frac{1}{N^{2}}\sum_{j\neq\ell}\log|\lambda_{j}-\lambda_{\ell}|\right]\right\}d^{n}\lambda,$ (1.4)
@@
where $z_{0}=z_{0}(t,\alpha)$ and $P_{3g-2}$ is a polynomial of degree 3g23g-2. This is a priori surprising since $z_{g}$, which may be used to count maps with an arbitrary number of vertices, only depends on $z_{0}$ (which is explicitly known from the string equation (1.12) below) and a finite number of parameters: the half-valence $\nu$ and the 3g13g-1 coefficients of $P_{3g-2}$. The connection with the generating functions $e_{g}$ was made in [Er14], which established that $e_{g}(t,\alpha)$ solves a Cauchy-Euler equation in $t$, with a forcing term that depends on the generating functions $\{z_{k}(t,\alpha),\ k\leq g\}$ and $\{e_{k}(t,\alpha),\ k<g\}$. Building on this result, [Er14] demonstrated that for $g\geq 2$, $e_{g}(t,\alpha)$ is also a rational function of $z_{0}=z_{0}(t,\alpha)$,
@@
where $Q_{5g-5-r}$ is a polynomial of degree 5g5r5g-5-r. This, which proved and extended a conjecture proposed for $\nu=2$ in [BIZ80], led to the partial fraction expansion
@@
These results made it possible to express the number of regular even-valent $g$-maps with $j$ vertices in terms of $z_{0}$ and a finite number of real coefficients (specifically 3g+j33g+j-3). Indeed, taking successive derivatives of (1.11) with respect to $t$ gives, for $j\geq 1$, [Er14]
@@ 2.1. Orthogonal polynomials, Painlevé equation, and center manifold
$\int_{\mathbb{R}}p_{n}(\lambda)p_{m}(\lambda)w(\lambda)d\lambda=\delta_{nm},\quad w(\lambda)=\exp\left(-N\left(\dfrac{1}{2}\lambda^{2}+\dfrac{r}{4}\lambda^{4}\right)\right).$
@@
$x_{1}=\dfrac{\mu_{2}}{\mu_{0}},\quad y_{1}=0,\quad\mu_{i}=\int_{-\infty}^{\infty}\lambda^{i}\exp\left(-N\left(\lambda^{2}/2+(r/4)\lambda^{4}\right)\right)d\lambda.$
@@
Since $x_{n}\to\infty$ as $n\to\infty$ for the Freud orbit, the corresponding sequence $\{u_{n}\}$ converges to 0. In addition since $x_{n}\sim y_{n}$ the sequence $\{s_{n}\}$ converges to $2$, and using (2.4) we see that the sequence $\{f_{n}\}$ converges to $2$ as well. In other words, the above transformation has the property that the Freud orbit converges to the fixed point $P_{\infty}$ with coordinates $(2,2,0)$ in the $(s,f,u)$ space.
@@ 2.3. The 4-valent case
Although the number, 3g+j33g+j-3, of possibly non-zero entries of $V^{(j)}$ increases with $j$, some of these coefficients may nevertheless vanish. As shown in [ELT23b], when $\nu=2$, the dynamics of the sequence $\{V^{(j)}\}$ simplifies: for $g\geq 2$, all of the potentially non-zero entries of $V^{(j)}$ are consecutively found in a band whose length, 5g55g-5, is independent of $j$. Consequently, the dynamics of the sequence $\{V^{(j)}\}$ can be recast into that of a finite-dimensional discrete dynamical system by writing
@@
where the $x_{k}^{(j)}$ represent the entries of $V^{(j)}$ that are possibly non-zero, and defining the (5g5)(5g-5)-dimensional map
@@ 4.1. Derivation of the mixed valence discrete dynamical systems
$\int_{\mathbb{R}}p_{n}(\lambda)p_{m}(\lambda)w(\lambda)d\lambda=\delta_{nm},\quad w(\lambda)=\exp\left(-NV_{\bf t}(\lambda)\right),$
@@ 4.2. Center manifold
Linearization about $P_{\infty}$ reveals that the five eigenvalues are $\lambda=1$ with multiplicity one, and $\lambda=-2\pm\sqrt{3}$, each with multiplicity two. Consequently, $P_{\infty}$ has a one-dimensional center direction. A center manifold expansion may be obtained by writing s,f,zs,\ f,\ z, and $w$ in powers of $u$, and solving order by order after demanding that the resulting curve be invariant under the dynamics of system (4.4) and tangent to the center direction at $P_{\infty}$. Such a calculation can easily be performed with the assistance of a computer algebra software. Using Maple [Map21], we obtain the following asymptotic expansion as $u\to 0$,
@@ 3 Existence
Mixed 3 & 4 In progress 4.2] In progress In progress
@@
Valence 4 $2\nu,\ \nu\in\mathbb{N}$ 3 Mixed 3 & 4
Results Related to   Map Counts $\bullet\ 0\leq g\leq 2$ [EMP08] $\bullet\ 0\leq g\leq 7$ (Recursive) [ELT23a] $\bullet\ 1\leq g\leq 7$ (Explicit) [ELT23b] $\bullet\ 0\leq g\leq 2$ [EMP08] $\bullet\ 2\leq g\quad$ (Hyper-   geometric functions) [ELT23b] $\bullet\ 0\leq g\leq 2\$[EP12] In progress
Previously Known Results Related to   Map Counts $\bullet\ 0\leq g\leq 2$ [BIZ80] $\bullet\ 0\leq g\leq 5$ [Pi06]    $\bullet\ 0\leq g\leq 5$ [DY17]    $\bullet\ g=3\qquad$ [BGM21] $\bullet\ 0\leq g\leq 2$ [BD13] $\bullet\ 0\leq g\leq 5$ [DY17]
@@ 5.1.2. Center manifold expansion
$\int_{\Gamma}p_{n}(\lambda)p_{m}(\lambda)w(\lambda)d\lambda=\delta_{nm},\quad w(\lambda)=\exp\left(-NP(\lambda)\right),$
@@ 5.2.1. An explicit identity involving the Gauss hypergeometric function
We have numerically verified the validity of Conjecture 5.1 for values of $\ell$ between 0 and 20, and values of $g$ and $j$ between 1 and 20. This conjecture was recently proven by M.L. Yattselev [Ya25].
@@ 5.2.3. Interlacing zeros of generating functions for 2-legged maps
where $S_{g-1}$ is a polynomial of degree g1g-1. In [ELT23a], we established such expressions for values of $g$ up to 7 and noticed that the zeros of the polynomials $S_{g-1}$, for $2\leq g\leq 7$, are all real and interlace. Further automation (with the help of a symbolic computation software) of the comparison of the genus and center manifold expansions performed in [ELT23a] will indicate whether this behavior continues for larger values of $g$. Using an alternate method to calculate the $z_{g}$’s, we were able to further verify this trend for $g$ up to 15. If this continues, understanding the structure of these polynomials, as well as the nature of the distribution of their zeros as $g\to\infty$, promises to be a rich and rewarding line of investigation.
@@ 5.2.4. Higher order asymptotic expansions for map counts
thereby clarifying the rate at which ${\mathcal{N}}_{2\nu,e}(g,j)$ converges to ${\mathcal{N}}_{2\nu,e}^{\infty}(g,j)$. Plots of the ratios ${\mathcal{N}}_{2\nu,e}(g,j)/{\mathcal{N}}_{2\nu,e}^{\infty}(g,j)$ are provided in [ELT23b] for $\nu=2$ and $2\leq g\leq 7$.
@@ 5.2.5. Map counts and Bernoulli numbers
We are grateful to M.L. Yattselev for pointing out that Equation (5.4) with $C^{(g)}={B_{2g}}/{4g(g-1)}$ is a special case of Gelfand’s reciprocity relation for Bernoulli numbers, thereby confirming the above conjecture. This relationship provides a novel method to calculate the orbifold Euler characteristic $\chi_{orb}(M_{g})$ from the generating functions $e_{g}(z_{0})$.